OpenStudy (anonymous):

Substitute the given values into the given formula and solve for the unknown variable. If necessary, round to one decimal place. Use the approximation 3.14 for pi. C=44

OpenStudy (anonymous):

c=circumference? c=2pi r or d*pi 44=d pi 44/pi=d

OpenStudy (anonymous):

yes c= circumference

OpenStudy (anonymous):

then d= 44/3.14 r= 22/3.14

OpenStudy (anonymous):

thank you...got the answer...